An alcoholic drink has 5.5% alcohol by volume. The label on a 375mL bottle says it contains 1.6 standard drinks.
a) How many mL of alcohol are in a 375 mL bottle?
b) It is recommended that a fully licensed male driver should have a maximum of one standard drink every hour. Express this as a rate in mL/minute, correct to 3 significant figures.

Answer by Shin123(626)

375*0.055=20.625 mL of alcohol. 375=(1.6)x 375/8=(0.2)x x=234.375/60=3.90625 or about 3.906 mL/min.